Two narrow slits emitting light in phase are separated by a distance of 1.0 cm. The wavelength of the light is 5.0 × 10-7 m. The interference pattern is observed on a screen placed at a distance of 1.0 m. (a) Find the separation between the consecutive maxima. Can you expect to distinguish between these maxima? (b) Find the separation between the sources which will give a separation of 1.0 mm between the consecutive maxima.

Grade:10

1 Answers

Jitender Pal
9 years ago
SOL. Given that, d = 1 cm = 10^–2 m, λ = 5 * 10^–7 m and D = 1 m a) Separation between two consecutive maxima is equal to fringe width.| So, β = λD/d = 5 * 10^-7 * 1/10^-2m = 5 * 10^-5 m = 0.05 mm. b) When, β = 1 mm = 10^-3 m 10^-3m = 5 * 10^-7 * 1/D ⇒ D = 5 * 10^-4 m = 0.50 mm.